5. At a given time a shaft is rotating 50 rpm about a fixed axis, 20 s later it is rotating 1050 rpm. Solve its angular acceleration
1
Expert's answer
2020-07-14T08:54:40-0400

"\omega_1=2\pi\nu_1=2\pi\cdot\frac{50}{60}=5.23rad/s"


"\omega_2=2\pi\nu_2=2\pi\cdot\frac{1050}{60}=109.9rad/s"


"\epsilon=\frac{\Delta\omega}{\Delta t}=\frac{109.9-5.23}{20}=5.23rad/s^2" Answer.